The demand curve for luminous socks is given by

Q=50 - 0.5P

and the total cost function for any curve in the industry is C=4Q

a) what is the marginal cost of the firm

b)what is the change in price for a one-unit increase in output?

c) if an industry wee perfectly competitive what would be the industry output and price? Explain the procedure used to determine industry output and price.
Answer #1

a. MC = d/dx(4Q) = 4

b. For a one unit increase in output, there is a decrease of $1 price

P

Q

0

50

2

49

4

48

c. Under perfect competition, the price is determined where MC=MR = 4

and the output at P = 4 is Q = 50-0.5P = 50-0.5*4 = 48